The longest side of a triangle is 4 more than three times the shortest side and the remaining side is 6.2cm. Find the lengths of each side if the perimeter is 23cm.
Answer by ankor@dixie-net.com(22740)

The longest side of a triangle is 4 more than three times the shortest side and the remaining side is 6.2cm.
Find the lengths of each side if the perimeter is 23cm
:
let s = the shortest side
then
(3s+4) = the longest side
:
A simple equation
(3s+4) + s + 6.2 = 23
4s + 10.2 = 23
Find s, then you can find the long side too